### v1

# Current validation practice undermines surgical AI development

Abstract.
Abstract: Surgical data science (SDS) is rapidly advancing, yet clinical adoption of artificial intelligence (AI) in surgery remains severely limited, with inadequate validation emerging as a key obstacle. In fact, existing validation practices often neglect the temporal and hierarchical structure of intraoperative videos, producing misleading, unstable, or clinically irrelevant results. In a pioneering, consensus-driven effort, we introduce the first comprehensive catalog of validation pitfalls in AI-based surgical video analysis that was derived from a multi-stage Delphi process with 91 international experts. The collected pitfalls span three categories: (1) data (e.g., incomplete annotation, spurious correlations), (2) metric selection and configuration (e.g., neglect of temporal stability, mismatch with clinical needs), and (3) aggregation and reporting (e.g., clinically uninformative aggregation, failure to account for frame dependencies in hierarchical data structures). A systematic review of surgical AI papers reveals that these pitfalls are widespread in current practice, with the majority of studies failing to account for temporal dynamics or hierarchical data structure, or relying on clinically uninformative metrics. Experiments on real surgical video datasets provide the first empirical evidence that ignoring temporal and hierarchical data structures can lead to drastic understatement of uncertainty, obscure critical failure modes, and even alter algorithm rankings. This work establishes a framework for the rigorous validation of surgical video analysis algorithms, providing a foundation for safe clinical translation, benchmarking, regulatory review, and future reporting standards in the field.
1. Main
The research field of surgical data science (SDS) was formally introduced in 2017, establishing it as a distinct field at the intersection of surgery, data science, and artificial intelligence (AI)111Throughout this manuscript, we use the term ”surgical” in a procedural sense, consistent with the definition of SDS as encompassing ”all clinical disciplines in which patient care requires intervention to manipulate anatomical structures with a diagnostic, prognostic, or therapeutic goal, such as surgery, interventional radiology, radiotherapy and interventional gastroenterology” [maier2017surgical]. This includes closely related interventional domains such as endoscopy, which share video-centric data, workflows, and validation challenges. [maier2017surgical]. The highly interdisciplinary domain leverages data acquisition, analysis, and modeling to enhance surgical decision-making, execution, training, and patient outcomes throughout the entire surgical care pathway. Since its beginnings, SDS has shown remarkable growth in AI-based publications [li2025artificial] (e.g., [hashimoto2019computer, madani2022artificial, kiyasseh2023vision, khan2024artificial]). However, clinical translation of SDS methods remains limited. For example, while over 1,000 AI-enabled medical devices have received authorization from the US Food and Drug Administration (FDA) since 2017, only six AI-enabled devices have been specifically approved by the FDA General and Plastic Surgery Devices panel, and only 17 for the field of gastroenterology-urology222https://www.fda.gov/medical-devices/software-medical-device-samd/artificial-intelligence-enabled-medical-devices as of 2025.
While multiple factors contribute to the limited clinical translation of SDS methods (e.g., workflow integration or regulatory hurdles), a key bottleneck lies in the lack of robust and rigorous validation of SDS algorithms as a prerequisite for safety and efficacy. Previous works have shown that validation in AI-driven medical image analysis is often flawed [carstens2025artificial, reinke2024understanding, funke2023metrics, kofler2023we, vaassen2020evaluation]. For example, it has been shown that the choice of metrics that do not reflect the underlying biomedical research question can severely undermine the validity of validation outcomes [reinke2021common, reinke2024understanding]. To address this problem, the Metrics Reloaded initiative introduced a comprehensive list of metric-related pitfalls as well as a problem-aware metric recommendation framework guiding researchers in finding the most suitable metrics for problems related to classification (image-level classification, object detection, semantic/instance segmentation) [maier2024metrics].
While this work has gained much support in the research community, it was designed for image-based problems. Unlike radiology or digital pathology, which are image-centered, AI-enabled optimization of intraoperative surgical behaviors is largely dependent on the analysis of spatiotemporal (video) data of the surgical field. Any SDS system that aims for clinical impact must therefore account for temporal dynamics. Fig. 1 exemplifies how neglecting temporal aspects in validation can lead to fundamental pitfalls, e.g.:
annotation inconsistencies across frames can result in misleading metric values, even if predictions are correct,
commonly used metrics may fail to capture temporal stability, rewarding flickering predictions with artificially high scores, and
simple, un-weighted aggregation over video frames may obscure poor performance during clinically critical phases.
Although it seems obvious that temporal relations should be taken into account, our findings reveal that this is rarely done in practice. This becomes particularly apparent during result aggregation, where videos are typically split into frames that are treated as independent images, thus ignoring temporal continuity and structural dependencies. However, surgical videos contain many temporally adjacent frames that are highly redundant and strongly correlated. Naïve aggregation across such frames violates the assumption of independent and identically distributed (i.i.d.) samples, which underlies many statistical analyses including confidence interval (CI) estimation and significance testing. As a result, the overall performance score can be biased and misleading, particularly when redundant frames dominate over clinically critical but less frequent moments.
The present work aims to initiate a paradigm shift in the validation of surgical AI algorithms. It evolved from an international, multidisciplinary effort initiated at a dedicated workshop during the 2023 annual meeting of the Society of American Gastrointestinal and Endoscopic Surgeons (SAGES) – one of the leading international societies for minimally invasive surgery. The workshop brought together leading experts in SDS, surgery, and AI, and laid the foundation for translating findings from the Metrics Reloaded initiative into the video-centric context of surgery. Building on the outcomes of this workshop, we subsequently launched a large-scale Delphi process that brought together a global panel of more than 90 experts to systematically identify, refine, and consolidate validation pitfalls specific to surgical video analysis.
Specifically, this work makes the following pioneering contributions to the safe clinical adoption of surgical AI:
Consensus-based catalog of validation pitfalls: We introduce the first comprehensive list of pitfalls in the validation of surgical AI, together with their potential consequences and real-world risks. This catalog resulted from a combined approach including a literature review, agentic internet research, and a consensus-driven expert process.
Evidence for high occurrence of pitfalls: Through a systematic literature review, we demonstrate that these pitfalls frequently occur in current surgical AI studies.
Experimental demonstration of impact of pitfalls: Using real surgical data, we experimentally quantify how these pitfalls distort performance assessment and mask critical failure modes.
2. Results
Over a period of about three years, we conducted a structured process involving a total of 91 experts from surgery, computer vision, and data science across 68 institutions, ensuring a broad range of perspectives across both surgical practice and technical disciplines. A core of the initiative was a hypothesis-generating workshop with dedicated focus group discussions at the annual SAGES 2023 meeting and subsequently evolved into a multi-stage Delphi process. This iterative process enabled the development of a comprehensive theoretical foundation and a consensus-based catalog of validation pitfalls related to surgical video analysis, supported by a systematic review demonstrating their prevalence and by experiments quantifying their impact on performance assessment.
2.1. A multi-stage, multi-stakeholder Delphi process revealed numerous pitfalls in surgical AI validation
To systematically identify common flaws in validating AI for surgical video analysis, we combined empirical evidence with structured expert consensus. Our method rested on three pillars: (1) a traditional literature review in PubMed and Google Scholar, (2) agentic internet search tools, and (3) a four-stage Delphi process involving SDS experts and clinicians (see Methods (Sec. 4) for further details). This combined approach enabled us to compile, refine, and validate a comprehensive list of pitfalls that may compromise surgical AI validation.
The resulting pitfall catalog is summarized in Fig. 2 and was structured into three key pitfall categories: [P1] pitfalls related to data, [P2] pitfalls related to metric selection and configuration, and [P3] pitfalls related to metric aggregation and reporting. Each category represents a distinct level at which pitfalls can compromise the validity of validation outcomes. Each pitfall was mapped to specific potential consequences, such as introduction of biases, unreliable performance assessment, or undetected failure modes, and to associated real-world risks (e.g., regulatory delay, compromised surgical safety), as illustrated in Fig. 2. Concrete definitions are provided in Tab. SN 2.1 (Supplementary Notes), with visual examples in Figs. 1, 4-6, and Extended Data Figs. 1- 9. Below, we outline the collected pitfalls identified within each category, highlighting why they are particularly critical in surgical AI validation. While some issues are known from other machine learning domains, the surgical setting adds unique layers of complexity and severity.
[P1] Pitfalls related to data
Flaws in how data are acquired, curated, or partitioned that can affect the validity of subsequent performance claims.
P1.1: Non-representativeness and low relevance of data: The lack of representative data is a familiar challenge in machine learning, where diversity is key for reliable validation. In surgery, however, this problem is amplified by diverse sources of variability, including procedures varying across hospitals, operating room (OR) setups, surgical technique, variations in data quality (Extended Data Fig. 1a), pronounced geographical imbalance of available datasets (Extended Data Fig. 1b), and even surgical team dynamics.
P1.2: Limited sample size/test cohorts, or data imbalance: Sample size limitations are widely discussed in medical imaging AI (e.g., [christodoulou2025false]), but surgical data is particularly hard to collect. Many surgical procedures are not routinely recorded, and rare but safety-critical events may appear only occasionally. Even single videos can last hours and require extensive annotation [carstens2025artificial]. Moreover, the resulting datasets often exhibit substantial class imbalance and small, heterogeneous test cohorts, which can lead to unstable and unreliable performance estimates (Extended Data Fig. 2).
P1.3: Existence of spurious correlations within data: Vision models often exploit accidental cues, but surgical datasets introduce additional confounders such as specific scopes, surgical team compositions, or OR layouts (Extended Data Fig. 3a).
P1.4: Incomplete annotation or missing contextual information: Incomplete labels reduce validity in general machine learning, but surgical videos heavily depend on temporal context, which is compromised by annotating only a fraction of frames (Extended Data Fig. 3b).
P1.5: Unreliable or inconsistent annotation: High inter-rater variability is a common problem in medical imaging AI [kolbinger2025appendix300, joskowicz2019inter, funke2023metrics], yet surgical tasks are especially ambiguous (e.g., phase transitions, fine tool-tissue interactions). Even trained raters frequently disagree and maintaining consistency over frames of a video of multiple hours is particularly challenging (Fig. 1a).
P1.6: Poor data splitting: Data leakage is a common problem in machine learning. However, surgical videos are highly redundant, i.e., a single patient (= case) may generate thousands of dependent frames. Without strict separation between data subsets, results may measure memorization rather than true generalization to unseen procedures (Extended Data Fig. 4).
[P2] Pitfalls related to metric selection and configuration
Flaws arising from the choice or setup of performance metrics that may distort results or fail to reflect clinically meaningful outcomes.
P2.1: Mismatch of metrics and clinical needs: Choosing metrics that reflect the actual clinical objectives is important for every research field [maier2024metrics]. However, the gap between existing metrics and needs is particularly wide for surgical applications. Systems must operate in real time, keep latency within safe limits, and produce outputs that remain temporally stable across rapidly changing scenes. Existing measures rarely capture those aspects (Extended Data Fig. 5), and, for many clinically relevant aspects, no established metric may yet exist.
P2.2: Lack of common metrics that assess temporal aspects: Temporal reasoning is crucial in surgery. However, standard metrics operate on a frame level or by simply aggregating over frames without considering temporal dynamics (Fig. 1b). This omission means that errors during safety-critical phases, or instability over time, may remain hidden.
P2.3: Inappropriate metric selection for handling annotation uncertainties: Ambiguous labels occur across domains, but phase boundaries or subtle tool-tissue contacts, among others, make ambiguity even more complex in surgical videos. In surgical reality, human spatiotemporal understanding is often associated with considerable inter-rater variability and inherent uncertainty [kolbinger2025appendix300]. Metrics assuming confident labels can either exaggerate or underestimate errors (Extended Data Fig. 6).
P2.4: Lack of metric robustness across varying real-world conditions: Metrics should be consistent across real-world conditions. In surgery, even well-defined measures may behave inconsistently when OR conditions vary – for example, when lighting changes, smoke, or blood partially obscure the field of view, the camera moves, or objects change in size or move in and out of view. Such factors can distort point estimates, despite stable model behaviors. (Extended Data Fig. 7).
P2.5: Non-standardized configuration and definition of metrics: In many AI applications, inconsistent thresholds or averaging rules reduce comparability. In surgery, even slight differences in how a metric is configured, such as overlap thresholds or smoothing windows, can obscure failures in short, safety-critical steps (e.g., vessel clipping) or make studies with the same metric incomparable (Extended Data Fig. 8a).
P2.6: Non-suitability of hyperparameters from unrelated domains: Translating hyperparameters from generic vision tasks is common practice to support standardization. However, in surgical videos, object sizes, motion speed, and safety requirements differ; often, a coarser threshold is already sufficient to track where instruments or anatomy are located within the scene, while overly strict settings may conceal whether an algorithm can follow events robustly over time (Extended Data Fig. 8b).
P2.7: Intrinsic limitations of individual metrics: Every metric comes with limitations. Translating standard metrics to surgical video analysis introduces additional challenges. Single scores may overlook brief but high-risk errors, fail to capture stability across time, or ignore how mistakes propagate through multi-step procedures.
[P3] Pitfalls related to metric aggregation and reporting
Flaws in summarizing and presenting results that can potentially lead to misinterpretation and undermine transparency.
P3.1: Non-independence within the test set: Correlated samples are a known concern in performance validation. In surgical video analysis, however, thousands of adjacent frames or several clips from the same patient may appear in the test set, inflating apparent confidence and masking how a system behaves on genuinely new procedures (Fig. 4).
P3.2: Clinically uninformative aggregation: Aggregating scores is common practice, but simple averaging over all frames can hide poor performance during high-risk phases (Fig. 1c) or overweight patients with longer procedures, obscuring performance on shorter, potentially riskier cases. The lack of stratification by clinically relevant conditions further conceals failure modes that may only appear under specific challenges or surgical contexts (Fig. 5).
P3.3: Lack of contextualization of performance values: Point estimates without context are problematic in any field, but even more so in surgical AI as they can be highly misleading (Extended Data Fig. 9a). For example, a high Accuracy may mainly reflect routine phases, while errors cluster in moments of adverse events.
P3.4: Lack of uncertainty reporting: Uncertainty estimates are often neglected in AI validation. For surgical systems, missing information on confidence or calibration limits the clinicians’ ability to decide when model outputs can be trusted during an operation (Extended Data Fig. 9b). This effect is even more critical if hierarchical data structures are not considered (see P3.1).
P3.5: Insufficient reporting: Sparse or incomplete reporting undermines reproducibility everywhere, yet for surgical applications, the consequences are immediate. Without clear descriptions of data sources, inclusion criteria, metric definitions, and aggregation methods (Fig. 6), it is impossible to judge whether results cover critical steps or rare complications.
2.2. Validation flaws are widespread in common practice
While pitfalls can theoretically occur in any validation study, their actual prevalence in state-of-the-art surgical AI publications remained unclear. To address this, we conducted a systematic screening of all papers at the 2023 Medical Image Computing and Computer Assisted Intervention (MICCAI) conference that applied deep learning methods to surgical data. As the leading international conference for medical image computing and computer-assisted interventions, MICCAI provides a representative overview of current practices in surgical AI. Key results of this analysis are summarized in Fig. 3 and in Suppl. Note 3.
Of all papers meeting the inclusion criteria (n = 46), 74% used surgical video data. The screening revealed several shortcomings across datasets, metrics, aggregation procedures, and reporting practices.
Surgical data cohorts were typically small and fragmented. The median dataset contained 37 training, 10 validation, and 22 test videos (minimum: 6 training, 2 validation, 2 test videos). 79% of datasets were only used once, with 38 distinct datasets identified across papers. Moreover, only 47% explicitly reported an untouched test set, while this was unclear in 37% of papers.
Temporal and modality-specific considerations were largely missing. A total of 77% of papers did not assess properties specific to temporal data, and only a single paper used a temporal consistency metric (see Extended Data Fig. 10).
Metric use was heterogeneous and rarely justified. Across all papers, 41 metrics were used only once (see Extended Data Fig. 10). The most commonly used metric was Accuracy. While only 30% properly justified their metric choice, 20% of those justified by popularity alone. In addition, in 80% of papers, it was unclear whether clinical relevance had been considered when selecting metrics.
Aggregation practices were insufficient. The aggregation procedure was unclear or not described at all in 66% of papers (see Fig. 6). Among studies involving hierarchical structures, (e.g., patient-level), only 5% explicitly accounted for their dependencies (see Fig. 4). Furthermore, 80% did not contextualize performance values, for example against human baseline or clinical thresholds.
Reporting was incomplete and rarely reproducible. 59% of papers did not (fully) report dataset sizes. Only one paper reported CIs, and one reported prediction intervals. Notably, 98% of papers did not report inter-rater variability. Ethical, legal, and social aspects (ELSA) were largely absent; 78% lacked ethical reporting, 89% ignored fairness or biases, and 91% omitted social, legal, or governance considerations. Ultimately, only a single paper reported sufficient detail to enable reproducibility; all others were missing relevant details in one or several aspects.
2.3. Experiments demonstrate consequences of pitfalls using real-world data
To move beyond theoretical examples, we experimentally investigated the consequences of selected pitfalls using representative surgical datasets. As surgical videos are typically long, temporally structured, and safety-critical, the manner in which results are aggregated and reported can strongly influence the visibility and interpretation of algorithm weaknesses. Given the video- and time-sensitive nature of surgical video data, and the limited empirical evidence in the literature, we focused our experiments on pitfalls concerning metric aggregation and reporting [P3]. All experimental procedures are described in the Methods (Sec. 4).
Dependent test samples inflate confidence
Surgical data are inherently hierarchical. Frames from a single video are not independent, as they come from the same patient and are influenced by factors such as the performing surgeon, hospital, or used surgical tools. Ignoring this structure can lead to unreliable performance estimates. As shown in the previous section, only a fraction of studies properly address hierarchical data, yet little empirical evidence exists on how this practice may affect uncertainty estimates.
To determine this impact, we analyzed two of the most widely used real-world datasets: (1) the Robust Medical Instrument Segmentation (RobustMIS) 2019 challenge dataset [ross2021comparative] for binary instrument segmentation and (2) the CholecTriplet dataset [nwoye2023cholectriplet2021] for surgical action triplet recognition. For both tasks, we compared CIs derived from a naïve bootstrap approach, which does not account for hierarchical dependencies, against CIs from a hierarchical bootstrap (Fig. 4).
Accounting for hierarchical data structure led to substantially wider CIs, which reflects the additional variance introduced at the video (i.e., patient) level that is ignored when assuming independence across all samples (naïve approach). Concretely, for binary segmentation, the CI widths increased by a median factor of more than 2 for both Dice similarity coefficient (DSC; 2.4x wider) and Normalized surface Dice (NSD; 2.2x wider). For surgical action recognition, CIs were 13.5x wider for mean Average Precision (mAP), 11x wider for weighted mAP, and 7.1x wider for top-5 Accuracy. These findings demonstrate that ignoring data dependencies can drastically understate model uncertainty, potentially giving a false sense of algorithm reliability in surgical settings.
Averages hide critical failures
Many studies in surgical AI summarize results as a single overall score, averaging performance across all frames or cases. While this practice is convenient, there is little evidence on how such aggregation may conceal errors under conditions where reliability is most critical for patient safety. In surgery, visual and technical challenges, such as smoke or rapid tool motion, can strongly affect algorithm robustness, yet these factors are rarely analyzed in validation reports.
To shed light on this problem, we compared global results with stratified analysis on multi-instance instrument segmentation results from the RobustMIS challenge [ross2021comparative], using metadata describing various relevant, potentially confounding image properties [ross2023beyond] (Fig. 5). While aggregated DSC scores suggested stable performance across algorithms, stratification by clinically relevant conditions revealed considerable performance drops. For instance, the median DSC decreased by 0.36 (up to 0.51 for one algorithm) in frames with intersecting instruments, and smaller but clear declines appeared for smoke and motion artefacts. Purely reporting globally aggregated values can therefore be highly misleading, whereas stratification exposes failure cases in safety-critical situations.
Aggregation choices can flip the winner
Surgical video analysis involves several hierarchy levels and the way results are aggregated across them can substantially affect reported performance. Yet, as shown in Sec. 2.2, the majority of studies do not explain the exact aggregation procedure, leaving readers unable to judge whether rankings or scores reflect clinically meaningful behavior. Here, we systematically assessed how different aggregation strategies influenced conclusions.
We analyzed results from the binary segmentation task of the RobustMIS challenge [ross2021comparative] and applied six different aggregation strategies: frame-wise, video-wise, phase-wise, phase-wise video-wise, video-wise phase-wise, and weighted phase-wise (see Methods (Sec. 4) for detailed descriptions of each strategy). We then compared the results for each strategy with the default frame-wise aggregation (Fig. 6). Similarly to the original challenge, we used the 5% percentile as the aggregation operator to reflect worst-case performance.
The median Kendall’s tau correlation [kendall1938new] across the different rankings compared to the default ranking was 0.68, indicating high variance in the leaderboards. The original winner changed in 80% of rankings, with a median absolute rank change of 1 and a maximum change of 3. Negative rank shifts occurred in 58% of cases, positive in 28%, and no change in 14%. As shown in Fig. 6b, when performance differences between algorithms were modest, even small changes in the aggregation led to substantial ranking shifts, questioning the reliability of the winner. These findings show that minor reporting omissions such as unspecified aggregation can substantially affect ranking conclusions and potentially influence which algorithms are prioritized for clinical translation.
3. Discussion
Our work provides the first comprehensive, expert-driven taxonomy of validation pitfalls in surgical AI, supported by empirical evidence and experimental results. By linking methodological flaws to surgical risks, our framework highlights the need for rigorous validation to ensure safe and effective AI deployment. Our findings demonstrate that common validation practices frequently ignore the temporal and hierarchical structure of surgical data, leading to overconfident or clinically irrelevant conclusions. Through a Delphi process with experts from surgery, machine learning, biostatistics, and regulatory affairs, we specifically identified and contextualized 18 critical pitfalls. This multi-stakeholder approach ensured that the collected pitfalls are both technically sound and clinically meaningful, making them actionable for a broad range of players, from algorithm developers to regulatory reviewers.
The collected pitfalls span all stages of the validation process. For example, at the data level, geographical imbalance of surgical and endoscopic datasets (Extended Data Fig. 1b) shows how limited geographic diversity can restrict representativeness, as many public benchmarks have historically been acquired from a narrow range of regions. Recent initiatives, such as the Critical View of Safety (CVS) Challenges 2024 and 2025 [alapatt2025sages], represent encouraging steps toward broader global inclusion, yet most benchmark datasets still predominantly reflect surgical practice patterns from Western regions. At the metric level, misalignment between chosen metrics and clinical objectives remains common, as metrics tailored to surgeon-specific needs or temporal aspects are often lacking. Consequently, studies frequently rely on simple frame-based scores such as Accuracy or DSC, which fail to capture the temporal and hierarchical complexity of surgical workflows and can obscure clinically relevant weaknesses. At the aggregation level, naïve frame-wise aggregation across temporally dependent data can mislead performance estimates, concealing critical failure modes and even altering algorithm rankings.
In addition to establishing this taxonomy, we collected empirical evidence confirming that the identified pitfalls are widespread across the surgical AI literature. Our systematic screening revealed frequent issues such as unstratified aggregation, lack of uncertainty reporting, unjustified and unsuitable metric selection, and poor documentation of validation procedures. Despite temporality being inherent to intraoperative surgical workflows and despite the existence of temporal metrics, these were rarely used in practice. Generally speaking, although surgical video AI tasks often involve complex, structured data such as multiple classes, temporal dependencies, and hierarchical structure, validation in common practice frequently remains limited to simple Accuracy metrics. As shown by our experiments, naïve frame-wise validation without consideration of temporal aspects or data hierarchy can mask critical failure modes and lead to a substantial underestimation of uncertainty. Together, these findings underscore the need for a paradigm shift in common surgical AI validation practice.
Our work comes with several limitations. While our taxonomy targets surgical video analysis, its generalizability to other temporally structured domains (e.g., cardiology) remains to be explored. Methodologically, although our Delphi expert consortium included more than 90 international experts and covered a broad range of expertises, it may not have captured the full diversity of surgical subfields and regulatory perspectives. Participation also varied across Delphi rounds, raising the potential for biases in the weighting and selection of pitfalls. In addition, our systematic literature review only covered 46 articles. However, although the sample size was limited, MICCAI is the leading conference for medical image analysis and computer-assisted interventions and can thus be considered representative for the field. It should further be noted that our experiments only tackled selected pitfalls. While pitfalls such as data leakage have been demonstrated in the broader machine learning community, we explicitly focused on aspects that are especially critical in surgery: aggregation under temporal and hierarchical structures. Other pitfalls are illustrated in smaller real-data analyses in Extended Data Figs. 1- 9. Finally, our present work does not yet provide concrete solutions for each pitfall. However, developing concrete recommendations will be the focus of future work of the consortium.
Several open research directions emerge from our presented pitfall taxonomy. At a conceptual level, future work should move beyond surrogate metrics toward validation that reflects clinical benefit and patient outcomes. This includes defining what constitutes sufficient performance within specific clinical contexts and establishing comparability across heterogeneous tasks. Methodologically, clear validation phases for SDS systems that integrate governance and stakeholder input, as well as standardized reporting, need to be defined and appropriately integrated into clinical trial design. Ensuring robustness through post-deployment monitoring, addressing catastrophic-failure risk and safety considerations, and enabling effective human-AI collaboration in the OR will be equally crucial. From a technical and adoption perspective, progress depends on harmonizing label ontologies and annotation protocols across datasets, facilitating validation for multimodal data while protecting patient privacy, and assessing behavioral consistency of AI models across samples and software versions to ensure stability after updates or retraining. Finally, embedding clinician priorities, workflow impact, and real-time safety mechanisms should become integral validation goals.
In summary, while Metrics Reloaded [maier2024metrics] provided metric recommendations for image-based validation, our new framework extends this foundation to pitfalls stemming from the temporal and hierarchical complexity of surgical video analysis. We envision this work as a catalyst for improved validation practice and future benchmarking efforts. By raising awareness of widespread pitfalls, we aim to inspire a paradigm shift toward more robust, interpretable, and clinically grounded validation pipelines. By systematically mapping validation pitfalls to their consequences, this work offers a structured foundation for integrating validation quality criteria into clinical trial design, regulatory review, and publication guidelines. Concretely, our findings may inform the development or refinement of reporting standards for medical video analysis, such as TRIPOD-AI [collins2021protocol], DECIDE-AI [vasey2022reporting], or future domain-specific extensions. Going forward, the consortium will focus on translating these pitfalls into surgery-specific metric and aggregation recommendations, further advancing the reliability and clinical readiness of surgical AI models.
4. Methods
4.1. Identification of validation pitfalls through a multi-stage Delphi process and complementary searches
The pitfalls presented in this work were derived through a combination of approaches, centered on a multi-stage, consensus-driven Delphi process conducted by an international, multidisciplinary panel of experts. A Delphi process is a structured consensus-building approach in which experts provide input individually – typically through questionnaires – followed by rounds of controlled feedback and refinement [brown1968delphi]. This methodology is widely recognized in medicine as a way to establish best practices in areas in which the available evidence is limited, inconsistent, or missing [nasa2021delphi].
Our Delphi panel initially included 60 international experts from the SDS initiative. To broaden the diversity of the expertise, the consortium was gradually expanded to 91 members across 68 institutions, reflecting both technical and clinical backgrounds. The expert panel was composed of 31% clinical, 74% technical, and 5% shared expertise. 12% of experts were from industry. The majority of experts were affiliated in Europe (69%; mostly Germany (35%) and United Kingdom (12%)) and North America (25%; mostly United States of America (20%)), followed by Asia (7%) and Africa (2%).
This initiative started in March 2023 with an initial scoping survey to identify the most critical problems in validating surgical AI and to capture use cases lacking suitable metrics or showing discrepancies between metrics and the clinical needs (participation rate: 33%). Building on the survey findings, we held an in-person kickoff workshop at the SAGES annual meeting in Montréal, Canada (41% in-person participation). This workshop refined the project scope and set priorities, laying the groundwork for the following Delphi rounds. Based on the workshop discussions, participants agreed to focus the initiative on surgical video understanding, reflecting shared priorities across clinical and technical stakeholders. Following the workshop, the core team performed targeted literature searches to compile candidate pitfalls. In parallel, a joint retreat involving members from three research groups at the German Cancer Research Center (DFKZ), National Center for Tumor Diseases (NCT), and University College London (UCL) provided an additional forum to critically discuss and refine preliminary pitfalls based on practical experience and interdisciplinary perspectives.
In addition, the core team performed a literature review to identify additional pitfalls. Concretely, we explored two databases, namely PubMed and GoogleScholar, as well as a general Google search, and used the following search string: (”surgical data science” OR ”surgical artificial intelligence” OR ”surgical AI” OR ”surgical scene understanding” OR ”surgical video analysis”) AND (”validation” OR ”evaluation” OR ”metric”) AND (”pitfall” OR ”limitation” OR ”caveat” OR ”drawback” OR ”shortcoming” OR ”weakness” OR ”flaw” OR ”disadvantage”). PubMed returned no relevant results, while Google Scholar yielded 704 hits and a general Google search 94 results (30 non-peer-reviewed). Although these searches yielded several appropriate items, no comprehensive or structured overview of validation pitfalls was identified.
Based on the results of the first survey, the workshop and retreat, and the internal literature review, the core team established a preliminary catalog of pitfalls, including their categorization, which served as the starting point for subsequent refinement through the Delphi process. To reduce blind spots, we complemented this with leading agentic internet research systems, including the deep research tools from OpenAI (o3-based), Google (Gemini 2.5 Pro-based), and Perplexity Pro, to help identify potentially overlooked pitfalls. All additional suggestions from these tools were validated by the expert consortium for relevance and correctness.
In total, we conducted four Delphi rounds (participation rates: 64%, 70%, 48%, 53%). Round 1 confirmed the overall project scope, while round 2 refined the pitfall categorization and pitfall list, identified missing pitfalls, and collected supporting references for the evidence of pitfalls. Round 3 focused on linking pitfalls to consequences and risks, and round 4 sought final consensus on the pitfall catalog and pitfall categorization (agreement: 98%), as well as optional feedback on figures, experiments (agreement 90%), and open research questions.
4.2. Systematic review for prevalence of pitfalls
From all MICCAI 2023 papers (n = 730), we identified all papers related to SDS (n = 51). From those, five articles were excluded because they did not deal with deep learning-related methods, therefore, several questions did not apply (n = 46). Each paper was screened by two independent screeners, with a total of twelve screeners. Afterwards, a third senior screener compared results and resolved conflicts. In total, three senior screeners joined this last step, with the papers divided among them. The screening covered more general aspects such as task or surgery type, but specifically focused on evidence for the identified pitfalls. In line with the Delphi process and after identifying the final list of pitfalls, a follow-up screening, following the same process, was conducted to ensure evidence for all pitfalls.
4.3. Experimental Design
Data
To ensure the general relevance of our findings, we based our experiments on two of the most widely used datasets in surgical video analysis [carstens2025artificial] that have been used in international challenges, are highly cited in the field, and cover two key tasks in surgical video analysis: instrument segmentation and action recognition.
The RobustMIS challenge 2019 [ross2021comparative] consists of videos from 30 laparoscopic colorectal surgeries, covering three surgery types, namely rectal resection, proctocolectomy, and sigmoid resection, with 10 videos, i.e., patients, per surgery type. For the challenge, the data from rectal resection and proctocolectomy surgeries were used for training and internal testing (stages 1 and 2), while sigmoid resection cases (stage 3) were reserved for assessing generalization to an unseen surgery type. For our experiments, we restricted the analysis to data from stage 3 (sigmoid resection). The challenge consisted of three tasks: binary segmentation, multi-instance detection, and multi-instance segmentation. For our experiments, we leveraged results for both segmentation tasks. For binary segmentation, a total of ten algorithms participated, for multi-instance segmentation, seven algorithms participated. The challenge metrics included the DSC, which measures the overlap between prediction and reference, and NSD, which assesses boundary accuracy [maier2024metrics] for binary segmentation and their multi-instance variants (MI_DSC and MI_NSD) for the multi-instance segmentation task. We had access to the frame-level metric scores submitted by all participating teams, which allowed us to analyze the impact of validation choices across a diverse set of real-world algorithms. All results were used in anonymized form to ensure confidentiality.
In addition to the segmentation masks, we utilized additional structured meta-annotations for the RobustMIS data, describing the presence of common visual artifacts or image characteristics [ross2023beyond]. These meta annotations indicate whether specific challenges such as blood, smoke, or motion blur are present per frame and instrument.
We further used the data and task setup of the CholecTriplet challenge [nwoye2023cholectriplet2021], which is based on 45 laparoscopic cholecystectomy videos (CholecT45). The task involves recognition of surgical action triplets, with annotations for 100 triplet classes, each defined by a combination of instrument, verb, and target. For our experiments, we used a Swin-Base Transformer trained using multitask learning, incorporating information on the instrument, verb, and target and soft-labels generated using a multi-teacher approach [yamlahi2023self]. Model validation was performed using 5-fold cross-validation, following the official CholecT45 setup.
Experiment 1: Dependent test samples inflate confidence
This experiment investigated how ignoring data dependencies in temporally structured surgical video data can lead to severely underestimated model uncertainty. To ensure relevance across both low-level and high-level prediction tasks, we focused on two widely used benchmark tasks, instrument segmentation (RobustMIS) and surgical action recognition (CholecT45). Both datasets exhibit a clear hierarchical structure, with multiple correlated frames per patient case.
For the binary segmentation task, we used the results of the ten algorithms of the binary segmentation task of the RobustMIS challenge. The same metrics as in the original challenge were applied, namely the DSC and NSD. In this dataset, one hierarchical level was considered, namely the patient (i.e., the video; n = 10).
For the surgical action triplet recognition task, results were derived from the CholecT45 dataset using one algorithm (Swin-Base Transformer; see above). We calculated the original mAP [maier2024metrics] metric as done in the original challenge. Given the class imbalance across 100 triplet classes, we additionally computed a class-weighted mAP, and included the top-5 Accuracy for completeness. Metrics were calculated for each cross-validation fold. Again, we considered the patient (i.e., the video; n = 45) as the relevant level of hierarchy.
For both tasks, CIs were estimated using two resampling strategies: the standard bootstrap and the hierarchical bootstrap. In the standard (naïve) bootstrap, we performed resampling with replacement of all frames of the entire test set 1,000 times without considering the hierarchical structure [efron1994introduction]. For each resample, we calculated the mean metric for each bootstrap sample, and obtained the empirical quantiles from the resulting bootstrap distribution to calculate 95% CIs. For the surgical action triplet recognition task, resampling was applied across all frame-level predictions within each cross-validation fold.
In contrast, the hierarchical bootstrap explicitly accounted for dependencies at each hierarchy level [saravanan2020application]. We first resampled the videos, i.e., the higher hierarchy level, followed by resampling the individual frames within each sampled video. The mean metric was then computed across all resampled frames and videos. This process was repeated 1,000 times, and the empirical quantiles of the resulting metric means were used to estimate the CIs. For the surgical triplet recognition task, this procedure was applied separately within each cross-validation fold, with metrics averaged per video and per fold before CI estimation.
For both CI methods, we calculated the CI widths for each algorithm and metric as well as the ratio between hierarchical and naïve CI widths.
Experiment 2: Averages hide critical failures
This experiment investigated whether global (non-stratified) aggregation of metric scores can conceal algorithm weaknesses under challenging image conditions. To enable stratified analysis across clinically relevant image characteristics, we focused on the RobustMIS dataset, for which we had access to structured metadata on visual artifacts or image properties [ross2023beyond]. While the original study [ross2023beyond] employed these annotations to analyze model robustness across visual conditions, our analysis focused on how global aggregation can obscure property-dependent performance differences that are critical for assessing validation reliability. Specifically, we analyzed the multi-instance segmentation task of the RobustMIS challenge, using the MI_DSC scores from the seven participating algorithms.
For each algorithm, the median MI_DSC across all frames was calculated as the baseline performance (non-stratified). The median was chosen instead of the mean to reduce sensitivity to outliers. Stratified performance was then calculated by restricting the analysis to frames containing individual artifact types. The following properties were considered: blood, reflections, smoke, motion, overexposed, underexposed, intersecting instruments, and low-artifact scenes (i.e., scenes with one or fewer annotated properties). Because frames could contain multiple artifacts simultaneously, these subsets were not mutually exclusive.
For each algorithm, we calculated the median MI_DSC for the full dataset (non-stratified baseline) as well as within each artifact-specific subset. We then computed the difference between the stratified and non-stratified medians per algorithm. To summarize performance changes across algorithms, we further computed the median of these algorithm-level scores per artifact type and reported the difference between the stratified and non-stratified aggregated medians.
To estimate the uncertainty of these differences, we applied hierarchical bootstrapping with 1,000 iterations, considering one hierarchy level, namely the patient (i.e., the video). For each bootstrap iteration, the difference in median MI_DSC between the stratified and the baseline conditions was computed, and CIs were derived from the empirical quantiles.
Experiment 3: Aggregation choices can flip the winner
This experiment investigated how different aggregation strategies affect algorithm rankings, given that aggregation schemes are rarely reported in practice. We focused on the data from the RobustMIS challenge, as we had access to frame-level performance scores from all participating algorithms, enabling systematic comparison across different aggregation strategies. Specifically, we used the DSC scores of the ten participants of the binary segmentation task to simulate alternative ranking outcomes.
Six different aggregation strategies were calculated:
Frame-wise aggregation: Performance was aggregated equally over all frames of all videos, irrespective of procedure or phase. This approach served as the default, as it reflects common practice and was also applied in the original challenge.
Video-wise aggregation: For each video, performance was aggregated equally over all frames, irrespective of the surgical phase. The resulting video-level scores were then combined into one final aggregate.
Phase-wise aggregation: For each surgical phase, performance was aggregated equally over all frames, irrespective of the video. The resulting phase-level scores were then combined into one final aggregate.
Phase-wise video-wise aggregation: Performance was first aggregated for each phase within each video, resulting in rankings per phase and video. Rankings were then aggregated per phase over those rankings, and finally aggregated into one final ranking.
Video-wise phase-wise aggregation: Performance was first aggregated for each phase within each video, resulting in rankings per phase and video. Rankings were then aggregated per video over those rankings, and finally aggregated into one final ranking.
Weighted phase-wise aggregation: Performance was first aggregated for each phase, and the resulting phase-level scores were then combined into a global score using clinician-defined weights reflecting the clinical relevance and complexity of each phase. For simplicity, we adopted a straightforward weight assignment strategy, in which phases deemed less relevant or complex were assigned a weight of 1 (phases 0, 4, 5, 6, 9, 12), intermediate phases were assigned 2 (phases 2 and 10), and highly relevant or complex phases received a weight of 3 (phases 1 and 8) [maier2021heidelberg].
Across all aggregation strategies, the aggregation operator can vary, e.g., it could be the mean, median, or other percentiles. In line with the original challenge, we used the 5th percentile to reflect worst-case performance. To assess the agreement between the default (frame-wise) ranking scheme and alternative aggregation strategies, we calculated Kendall’s tau [kendall1938new], a rank correlation coefficient, quantifying the similarity between two orderings (Kendall’s tau = 1: perfect agreement, Kendall’s tau = 0: no association, Kendall’s tau = -1: complete disagreement).
5. Extended Data
6. Acknowledgements
A.R.: This work was initiated by the Helmholtz Association of German Research Centers in the scope of the Helmholtz Imaging Incubator (HI). L.M.-H.: This project was supported by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (NEURAL SPICING, 101002198), the National Center for Tumor Diseases (NCT), Heidelberg’s Surgical Oncology Program, the German Cancer Research Center (DKFZ). E.C.: This publication was further supported through state funds approved by the State Parliament of Baden-Württemberg for the Innovation Campus Health + Life Science Alliance Heidelberg Mannheim. O.C.: The research leading to these results has received funding from the French government under management of Agence Nationale de la Recherche as part of the “France 2030” program (reference ANR-23-IACL-0008, project PRAIRIE-PSAI), as part of the ”Investissements d’avenir” program (reference ANR-19-P3IA-0001, project PRAIRIE 3IA Institute and reference ANR-10-IAIHU-06, project Agence Nationale de la Recherche-10-IA Institut Hospitalo-Universitaire-6) and from the European Union’s Horizon Europe Framework Programme (grant number 101136607, project CLARA). R.D.: This project has received funding from the European Union’s Horizon Europe research and innovation programme under grant agreement No 101092646. D.D. is funded by NIH K23 EB034110. Q.D.: This work described in this paper was partially supported by a grant from the ANR/RGC Joint Research Scheme sponsored by the Research Grants Council of the Hong Kong Special Administrative Region, China and the French National Research Agency (Project No. A-CUHK402/23). G.F. is supported by Canada Research Chair in Computer-Assisted Surgery. S.Gia. is supported by the Royal Society URF R 201014. T.H. is a Consolidator Researcher, receiving financial support from the Distinguished Researcher program of Óbuda University. His work has been partially supported by ACMIT (Austrian Center for Medical Innovation and Technology), which is funded within the scope of the COMET (Competence Centers for Excellent Technologies) program of the Austrian Government. D.A.H. is supported by a grant from the American Surgical Association Foundation. F.R.K. receives support from the German Cancer Research Center (CoBot 2.0), the Joachim Herz Foundation (Add-On Fellowship for Interdisciplinary Life Science), the Central Indiana Corporate Partnership AnalytiXIN Initiative, the Evan and Sue Ann Werling Pancreatic Cancer Research Fund, and the Indiana Clinical and Translational Sciences Institute (EPAR4157) funded, in part, by Grant Number UM1TR004402 from the National Institutes of Health, National Center for Advancing Translational Sciences, Clinical and Translational Sciences Award. The content is solely the responsibility of the authors and does not necessarily represent the official views of the National Institutes of Health. J.L.L. received funding by the Swiss National Science Foundation (P5R5PM 21766), Novartis Foundation for medical-biological Research (#23C162), and the Vontobel Foundation (0867/2024). H.J.M. is supported by the NIHR UCLH/UCL Biomedical Research Centre. N.P.: This work has received funding from the European Union (ERC, CompSURG, 101088553, PI: Nicolas Padoy). Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. This work has also been supported by French state funds managed within the Plan Investissements d’Avenir by the ANR under reference ANR- 10-IAHU-02 (IHU Strasbourg). H.R.: Hong Kong Research Grants Council Collaborative Research Fund under Grant CRF-C4026-21G. S.S. is supported by the German Research Foundation (DFG, Deutsche Forschungsgemeinschaft) as part of Germany’s Excellence Strategy – EXC 2050/1 – Project ID 390696704 – Cluster of Excellence ”Centre for Tactile Internet with Human-in-the-Loop” (CeTI) of Dresden University of Technology. D.S. is supported by the Royal Academy of Engineering Chair in Emerging Technologies. M.W.: This work has been funded by the German Research Foundation (DFG, Deutsche Forschungsgemeinschaft) as part of Germany’s Excellence Strategy—EXC 2050/1—Project ID 390696704—Cluster of Excellence ”Centre for Tactile Internet with Human-in-the-Loop” (CeTI) of TUD Dresden University of Technology and by the German Federal Ministry of Research, Technology, and Space within the “Surgical AI Hub Germany” project (grant number BMBF 02K223A110).
We would like to thank Johannes Bender, Patrick Beyersdorffer, Isabel Funke, Susu Hu, Alexander Jenke, Denise Junger, Piotr Kalinowsky, Lucas Luttner, Keno März, Leon Mayer, Dominik Michael, Wenyao Xi, Jinjing Xu, and Mona Zeinodin, for fruitful discussions on the framework.
We would like to thank Kevin Cleary and Ajit Sachdeva for participating in the SAGES 2023 workshop.
7. Conflicts of Interest
M.S.A. is a speaker of Medtronic and advisor of Johnson & Johnson. D.A. is a shareholder and has a leadership role in Scialytics. N.W.C.: Nil. O.C. reports having received consulting fees from Therapanacea (2022-2024). OC reports that other principal investigators affiliated to the team which he co-leads have received grants (paid to the institution) from Sanofi (2020-2022) and Biogen (2022-2023). OC reports that his spouse was an employee of myBrainTechnologies (2015-2023) and is an employee of DiamPark. G.F. declares the following conflict of interest: Johns Hopkins University, Harvard Brigham and Women’s Hospital. S.G. declares the following competing financial interests: he has or has had consulting relationships with Una Health GmbH, Lindus Health Ltd., Flo Ltd, ICURA ApS, Rock Health Inc., Thymia Ltd., FORUM Institut für Management GmbH, High-Tech Gründerfonds Management GmbH, DG SANTE, Prova Health Ltd, haleon plc and Ada Health GmbH and holds share options in Ada Health GmbH. T.P.G. is the Founder of SST. D.A.H. is a consultant for Medtronic. A.J. is an employee of Intuitive. F.R.K. declares advisory roles for Radical Healthcare, USA; and the Surgical Data Science Collective, USA, and has received research funding from Novartis. H.J.M. is employed by and holds shares in Panda Surgical Limited. P.M. is a co-founder and shareholder of Scialytics. H.N. is affiliated with Proximie Ltd, London, UK. N.P. is co-founder and own shares in Scialytics SAS. N.R. is an employee at NVIDIA. M.M.R. is now working at SAP Fioneer. D.S.: Medtronic, Odin Vision, Panda Surgical, EnAcuity, Helico Medical, Uncovr, Vope Medical. M.W. has received honoraria by KARL STORZ SE & Co KG.
Supplementary Notes
Contents
| SUPPL. NOTE 1 | Pitfall descriptions and literature evidence | SUPPL. NOTE 1 |
| SUPPL. NOTE 2 | Descriptions of consequences and real-world risks | SUPPL. NOTE 2 |
| SUPPL. NOTE 3 | Results of the systematic review per pitfall | SUPPL. NOTE 3 |
Appendix SUPPL. NOTE 1 Pitfall descriptions and literature evidence
To systematically identify common flaws in validating artificial intelligence (AI) for surgical video analysis, we performed a traditional literature review, utilized agentic internet search tools, and conducted a four-stage Delphi process involving surgical data science experts and clinicians. This resulted in a catalog of 18 pitfalls across three pitfall categories. Tab. SN 1.1 provides detailed descriptions of each pitfall, evidence from literature including both papers including these flaws and papers discussing them, as well as references to illustrations in the main manuscript.
| [P1] Pitfalls related to data | ||||
| ID | Pitfall | Description | Evidence | Illustration |
| P1.1 | Non-representativeness, low quality, and low relevance of data | The data used for training and/or testing does not accurately reflect the intended real-world use case or lacks clinical relevance. This may, for example, result from geographical imbalance, a lack of diversity in patient demographics, regional differences in surgical standards, or limited or outdated variations in surgical techniques. Additionally, data selection criteria may exclude challenging but clinically relevant cases, such as poor camera quality, low lighting conditions, challenging procedures, or imaging artifacts, which are crucial for assessing model robustness in realistic scenarios. The inclusion of irrelevant data, such as out-of-body frames or unintentionally recorded segments, may also dilute the training signal or introduce misleading patterns. The use of simulated or experimental data may further reduce real-world applicability, while proxy data can misalign validation with clinical objectives. Variability in video quality and preprocessing protocols across institutions may further compromise data fidelity and model reproducibility. | [godau2023deployment, kolbinger2024strategies, koch2024distribution, sellner2023semantic, kilim2022physical, castro2020causality, subasri2025detecting, lavanchy2024challenges, kassem2022federated, mascagni2022multicentric, kong2021accurate, junger2022state, bar2020impact, paranjape2023cross, engelhardt2018improving, sahu2021simulation, huaulme2023peg, kolbinger2023anatomy, kolbinger2024artificial, ding2022carts, childers2021same, jogan2024quality] | Extended Data Fig. 1 |
| P1.2 | Limited sample size/test cohorts, or data imbalance | The test data may be limited due to small sample size and/or the absence of validation on independent datasets, including those from different institutions, time periods, or populations (external validation), or from real-world clinical use after deployment (post-deployment validation). Limited data can mislead model development (e.g., algorithm selection), increase the risk of overfitting, or yield unreliable model performance validation, making it difficult to draw meaningful conclusions about the model’s real-world performance. These issues are amplified in the presence of class imbalance, where rare classes may appear in only one subset or fold, or be entirely absent, leading to inconsistent model performance depending on the specific validation or test set used. | [bar2020impact, kostiuchik2024surgical, yamlahi2023self, demir2023deep, killeen2023pelphix, maqbool2020m2caiseg, cheng2023deep] | Extended Data Fig. 2 |
| P1.3 | Existence of spurious correlations within data | Spurious correlations are statistical patterns in the data that do not reflect robust or generalizable associations with the target task. They can arise from biases in data collection, labeling, or contextual factors unrelated to the underlying clinical objective, and often reflect dataset-specific artifacts that fail to generalize. This may lead to short-cut learning, where a model exploits irrelevant patterns, for example, achieving seemingly good performance by associating certain surgical outcomes with specific recording conditions, such as particular camera types used at different hospitals, or inferring the presence of instruments from irrelevant features such as glove color, if such features are biased within the dataset. | [yuan2024advancing, banerjee2024bias, mahmood2021detecting, geirhos2020shortcut, dammu2023detecting, bykov2023finding] | Extended Data Fig. 3a |
| P1.4 | Incomplete annotation or missing contextual information | Annotations of the data are – intentionally or unintentionally – incomplete, leading to gaps in the information available for reliable validation. This may include missing metadata, partially annotated video sequences, the omission of specific objects, events, or surgical phases, or insufficient detail for the intended task. Missing metadata, such as age, for example, may lead to undetected confounders and thus biased models. Partially annotated video data may render the application of temporal metrics (e.g., for assessing continuity or stability across frames) infeasible and thus limit the ability to assess performance over time, such as continuity or stability across video frames. While some approaches may use partial annotations intentionally (e.g., in weak supervision), unmanaged or undocumented incompleteness may still compromise reliability and validation validity. The omission of relevant entities can result in inaccurate performance estimates, for example, underestimating sensitivity when relevant instances are not labeled, and may give a misleading impression of model accuracy and clinical utility. | [pangal2021guide, ward2021challenges, tajbakhsh2020embracing, twinanda2016endonet, funke2019video, peng2024reducing, nyangoh2023systematic, murali2023endoscapes] | Extended Data Fig. 3b |
| P1.5 | Unreliable or inconsistent annotation | Unreliable or inconsistent annotations can undermine the validity of model validation by introducing ambiguity, bias, or uncontrolled variability. This may result from the lack or insufficiency of a standardized annotation protocol, high inter- or intra-rater variability, or inconsistencies in labeling the same entity across frames. Annotation variability may also arise from annotator fatigue, differences in clinical expertise, or a lack of annotation training and auditing procedures. Additionally, annotation quality may depend on task complexity, which may require multi-rater annotation and inter-/intra-rater agreement analysis. Even if annotations are consistent, they may still be unreliable, for example, if they are based on weak reference sources such as unreliable sensors or flawed clinical definitions. These inconsistencies and sources of unreliability can lead to models learning ambiguous or conflicting patterns, reduce the reliability of reported performance metrics, and make it difficult to interpret model failures or compare results across datasets or studies. | [nwoye2023cholectrack20, pangal2021guide, twinanda2016endonet, sharan2023mvhota, cartucho2024surgt, bano2024placental, mao2024pitsurgrt, radsch2023labelling, luca2022impact, demir2023deep, zisimopoulos2018deepphase, ghamsarian2023cataract] | Fig. 1a |
| P1.6 | Poor data splitting | Data are not adequately split into training, validation, and test sets, leading to issues such as overfitting or data leakage. This could, for example, occur due to improper use of test splits for validation or model selection, non-stratified sampling, use of test splits for ablation studies, or the complete lack of a test set. A particularly harmful form of leakage can occur when data from the same patient is included in both training and test sets. This allows the model to leverage patient-specific characteristics, leading to overly optimistic performance estimates and reduced generalizability to new patients. The problem of data leakage is currently becoming even more severe with the emergence of generalist models trained on unknown cohorts, where the lack of transparency in the data composition makes proper separation and assessment especially challenging. | [funke2023metrics, liu2023skit, gao2021trans, wei2022boxpolyp, ahmad2024wavemamba, cervantes2021automatic, he2022empirical, nwoye2022data, kostiuchik2024surgical, jiang2021risk, abdulbaki2018surgical] | Extended Data Fig. 4 |
| [P2] Pitfalls related to metric selection and configuration | ||||
| ID | Pitfall | Description | Evidence | Illustration |
| P2.1 | Mismatch of metrics and clinical needs | Metrics are used that do not reflect the underlying needs of the clinical use case or fail to capture the impact of the model’s performance on clinical decision-making. This often reflects a focus on technical performance rather than clinical relevance. For example, a high frame-wise accuracy in surgical phase recognition tasks may not be clinically meaningful if the model frequently misidentifies short but clinically relevant phases or critical transitions between surgical phases. Another example includes specific operational requirements such as real-time capabilities, which are often not considered in performance validation, despite being critical for clinical usability. Moreover, clinical needs are shaped not only by the task itself, but also by the intended role of the AI system, for example, whether it supports decision-making or acts autonomously. | [kankanamge2025artificial, hashimoto2024foundation, reinke2024understanding, kolbinger2024strategies, brand2022frame, bar2020impact] | Extended Data Fig. 5 |
| P2.2 | Lack of metrics that assess temporal aspects | Metrics are used that do not assess algorithm properties specific to temporal data. For example, in surgical instrument segmentation, validating performance solely based on frame-wise Dice Similarity Score (DSC) may result in overlooking temporal inconsistencies, such as abrupt appearance or disappearance of instruments between consecutive frames. In addition to temporal coherence of predictions, metrics for evaluating real-time system behavior, such as latency or throughput, are often missing, despite their relevance in time-critical clinical settings. | [dergachyova2016automatic, boers2020improving, zhang2022retrieval, berlet2022surgical, sharan2023mvhota, luiten2021hota, kopf2021robust, funke2023metrics, brand2022frame] | Fig. 1b |
| P2.3 | Inappropriate metric selection for handling annotation uncertainties | Annotation uncertainty is not adequately addressed in metric selection, which can result in misleading performance validation. This uncertainty may arise from inter- and intra-rater variability in subjective tasks, low visibility in surgical videos, or inconsistent labeling of ambiguous regions (e.g., tissue boundaries or transition moments) over time. For example, ambiguity in defining event boundaries - such as the exact moment a surgical phase transition occurs - may introduce inconsistencies in annotations due to inter- or intra-rater variability and missing exact ground truth. | [dergachyova2016automatic, funke2023metrics, berniker2021probabilistic, berlet2022surgical, nwoye2022data] | Extended Data Fig. 6 |
| P2.4 | Lack of metric robustness across varying real-world conditions | The selected metrics lack robustness with respect to various real-world conditions such as data quality, acquisition settings, or clinical variability. For example, performance validation may be overly sensitive to changes in frame rate, image resolution, zooming, differences in surgical technique or intraoperative conditions, or annotation granularity. This includes common intraoperative phenomena such as smoke, motion blur, objects temporarily leaving the field of view, or changes in object size due to camera movement or zoom. These factors can distort metric behavior even when model predictions remain stable, and may substantially affect metric stability and reproducibility. | [nwoye2025surgitrack, czempiel2021opera, knoche2021susceptibility] | Extended Data Fig. 7 |
| P2.5 | Non-standardized configuration and definition of metrics | Metrics without standardized configuration within a specific use case or definition are used, often without providing details on the concrete formula. For example, hyperparameters such as thresholds for surgical instrument detection, temporal tolerance ranges for transitions in phase recognition (e.g., how many frames of deviation are accepted as correct), or weighting of different error types may not be clearly defined or reported. In addition, even when the same metric is used, differences in spatial or temporal application, such as calculating accuracy only during annotated segments versus the full video, or validating phase recognition at different frame rates, can lead to non-comparable results and misinterpretation. | [funke2023metrics, gao2021trans, jin2021temporal, jin2017sv] | Extended Data Fig. 8a |
| P2.6 | Non-suitability of hyperparameters from other domains | Hyperparameters or default metric configurations adopted from other domains (e.g., general computer vision) may not align with the specific underlying clinical needs. For example, in object detection, a high Intersection over Union (IoU) threshold may be appropriate in general computer vision tasks, but in surgical applications, lower IoU thresholds might be more suitable when only rough localization of objects (e.g., polyps) is needed [tran2023sources]. Similarly, default hyperparameters for event detection in action recognition may not account for the variability in surgical workflows. In marker-less tool tracking (e.g., [hein2025next]), for example, thresholds that define acceptable accuracy in everyday scenarios (e.g., a few millimeters) may be insufficient in surgical contexts, where sub-millimeter precision can be clinically critical. | [tran2023sources, ramesh2023dissecting, hein2025next] | Extended Data Fig. 8b |
| P2.7 | Intrinsic limitations of individual metrics | The used metrics harbor pitfalls related to their individual mathematical properties. Even if metrics are well-aligned with the clinical task, they may exhibit problematic behaviors such as sensitivity to class imbalance, non-linearity, or lack of interpretability. The suitability of any given metric should thus be analyzed in light of its known limitations. For example, accuracy may produce misleading values in highly imbalanced data sets. | [reinke2021common, reinke2024understanding] | |
| [P3] Pitfalls related to metric aggregation and reporting | ||||
| ID | Pitfall | Description | Evidence | Illustration |
| P3.1 | Non-independence within the test set | Failure to account for non-independence within the test set can yield misleading conclusions. Non-independence can occur when multiple frames from the same surgical video are included, or when data from the same patient, procedure, or institution are used. Aggregation, statistical analysis and reporting need to account for this lack of independence rather than assuming independent samples. For example, if frames from the same patient are used for validation, performance metric values may appear artificially high due to strong temporal correlation. Other possible consequences include biased statistical estimates, such as underestimated uncertainty or distorted means. | [kostiuchik2024surgical, funke2023metrics, kirtac2022surgical] | Fig. 4 |
| P3.2 | Clinically uninformative aggregation | Performance metrics are often simply aggregated, without accounting for the varying importance of different surgical phases or time segments, or for performance differences across clinically relevant conditions. For example, averaging over all frames may obscure poor performance during critical moments, and reporting unstratified metrics may hide failures in challenging scenarios such as for rare phases, low-quality recordings, presence of artifacts, or small anatomical structures, potentially diluting errors during rare but clinically significant events. Aggregating over meaningful temporal units and stratifying results by clinically relevant factors, including those affected by long-tail distributions (e.g., rare events or structures), as well as subgroup characteristics such as disease severity, surgical indication, or operator experience, enables more informative and clinically useful validation. Stratification can also improve transparency and fairness by highlighting differences in performance across subgroups. However, care must be taken in how metrics are applied and aggregated in stratified settings, as some metrics may behave non-intuitively when applied to imbalanced or small subgroups. | [funke2023metrics, kostiuchik2024surgical, zohar2020accurate, czempiel2023symphony, ross2023beyond, nwoye2023cholectrack20, nwoye2025surgitrack, huaulme2024global, hung2019experts] | Extended Data Fig. 9a |
| P3.3 | Lack of contextualization of performance values | Performance values are reported without sufficient context, making it difficult to assess their clinical or practical relevance. For example, results may be presented without comparison to human performance, inter-rater agreement, or a meaningful performance threshold that defines an acceptable error rate for the clinical task. Additionally, the significance of observed differences between models may not be validated, resulting in misleading interpretations. Without such contextualization, the clinical utility of a method may remain unclear. Additionally, aspects of how performance information is communicated to end users, including the design of visualizations, thresholds, and labels, can influence clinical perception and decision-making, and should be considered in the contextualization process. | [reinke2024understanding, jogan2024quality, sylolypavan2023impact, ward2021challenges] | Extended Data Fig. 9b |
| P3.4 | Lack of uncertainty analyses | Performance values are reported without conveying the uncertainty associated with the results, leading to overconfidence in the model’s performance. For example, confidence intervals, standard error, or standard deviations may be missing, concealing how much variability is present in the reported metrics, including across time or clinically relevant subgroups. | [christodoulou2024confidence, ban2021aggregating, shi2022attention, hao2023act, wang2023cascade] | Fig. 5 |
| P3.5 | Insufficient reporting | Reporting of results lacks sufficient detail and transparency, making it difficult to interpret, compare, and reproduce findings. Critical aspects, such as the strategy and rationale for aggregating performance metrics, details of metric computation, or the limitations of the validation data, are often underreported or unclear. In some cases, reporting may be selective rather than merely incomplete, which can lead to overestimation of performance and misleading impressions of clinical safety or utility. Additionally, established reporting guidelines tailored to specific purposes (e.g., CONSORT-AI for clinical trials, CLAIM for medical imaging) are frequently not followed, leading to incomplete or inconsistent documentation. Another common gap is the lack of documentation regarding data modifications introduced by acquisition hardware or manufacturer-specific processing pipelines (e.g., compression, interlacing, or automatic enhancement), which can affect model performance in subtle and unquantifiable ways. | [koccak2025adherence, loftus2023artificial, navarro2023systematic, marwaha2023appraising, junger2022state, funke2023metrics] | Fig. 6 |
Appendix SUPPL. NOTE 2 Descriptions of consequences and real-world risks
Each of the identified pitfalls was mapped to specific potential consequences, such as introduction of biases, unreliable performance assessment, or undetected failure modes, and to associated real-world risks (e.g., regulatory delay, compromised surgical safety). Tab. SN 2.1 and Tab. SN 2.2 provide detailed descriptions of consequences and risks.
| Consequence | Description |
| Introduction of biases | Biases in surgical AI refer to systematic errors arising from flaws in the data used for model development or validation, including imbalanced patient cohorts, underrepresentation of relevant scenarios, sampling artifacts, or spurious correlations. These flaws can lead to issues like shortcut learning, where models exploit statistically predictive but clinically irrelevant patterns. As a result, models may produce non-representative performance estimates, behave unreliably in real-world settings, and fail to generalize across diverse patient populations or surgical workflows. Such biases can not only reduce clinical trust but also risk perpetuating systemic blind spots in algorithmic behavior. |
| Unreliable performance assessment | Unreliable performance assessment refers to the generation of performance estimates that are unstable, inconsistent, or sensitive to uncontrolled aspects of the validation setup, or simply uninformative. This may result from non-robust experimental design, poor data splits, or metric configurations that are highly sensitive to implementation details. As a consequence, reported results may fluctuate across settings, making it difficult to draw reliable conclusions, compare methods, or understand true model behavior. |
| Insufficient transparency and/or reproducibility in performance assessment | Insufficient transparency and/or reproducibility in performance assessment refers to situations where results cannot be fully interpreted, contextualized, or independently reproduced. This may occur when critical details of the validation process are missing, unclear, or inconsistently applied, including aspects such as aggregation details or data preprocessing steps. As a result, others may be unable to replicate findings, understand the sources of variation in reported results, or assess their applicability to related clinical settings. |
| Undetected failure modes | Failure modes are patterns of incorrect or unsafe behavior that an AI model can exhibit under certain conditions. In surgical AI, these often arise from corner/edge cases – rare but clinically important scenarios in surgical video analysis that differ from typical training data – or from inadequate assessment of temporal behavior. Failing to properly validate these cases can lead to unreliable model behavior. When development datasets fail to capture such edge cases, these failure modes remain unnoticed during validation and may only surface in real-world use. Examples include unexpected anatomical variations, poor lighting conditions, occlusions, rare surgical complications, detection delay or instability of predictions across consecutive frames. Overemphasis on common cases leaves model performance in these critical situations unknown. |
| Leakage and/or model overfitting | Leakage and model overfitting result from improper use of data during model development and may lead to misleadingly high performance estimates that do not reflect how the model will perform in real-world clinical scenarios, reducing real-world applicability. Data leakage occurs when information from the data used for testing is also present during development e.g., through shared pre-processing, improper data splitting, or data re-use, violating the assumption that test data are independent and unseen. Model overfitting refers to selecting or optimizing models based on patterns that do not generalize beyond the data used during development. It often occurs when no properly untouched test set is reserved, allowing models to perform well in the corresponding test set only rather than showing true generalizability. |
| Suboptimal resource use | Suboptimal resource use refers to an inefficient use of time, funding, and (computational) resources in surgical AI development. This may occur when validation does not sufficiently reflect clinical needs, leading researchers to invest in models that address limited or misaligned problems and are unlikely to make a real-world impact. While such models may still contribute to methodological advancement, a lack of alignment with clinical priorities can limit the practical value of the work. Consequently, scientific progress may be slowed. |
| Real-world risk | Description |
| Overestimation of algorithm performance | Overestimation of algorithm performance refers to a mismatch between how well a surgical AI model is perceived to perform and how it actually behaves in clinical reality. This risk may arise from flawed or incomplete validation, misaligned performance metrics, or misinterpretation of reported results. As a consequence, models may appear more reliable, generalizable, or clinically useful than they truly are. Overestimation can lead to premature deployment, overreliance by clinicians, or inadequate oversight, increasing the risk of downstream errors, inefficiencies, or patient harm. |
| Unfairness of algorithms | Unfairness of algorithms refers to systematic performance differences across demographic groups, such as patients of different sexes, races, ethnicities, or socioeconomic backgrounds. This unfairness often arises from imbalanced training data, biased annotations, or model design choices that fail to ensure equitable performance. As a result, some groups may consistently receive more accurate predictions than others, leading to unequal treatment and raising significant ethical and clinical concerns. |
| Poor generalization to real-world applications | Poor generalization refers to the inability of an algorithm to maintain consistent performance across diverse real-world settings.This often results from shifts in context, technology, or clinical implementation that are not adequately captured during model development and validation. While a model may perform well in controlled settings, it may struggle when applied to different surgical procedures and teams, clinical workflows and sites, or medical devices (among others). For example, a model trained on data from one hospital may not generalize to another due to variations in equipment, imaging quality, or surgeon-specific practices. |
| Limited clinical utility | Limited clinical utility refers to the inability of algorithms to provide meaningful benefit in real-world clinical practice. This may occur when models fail to support clinical decision-making, integrate into workflows, or deliver reliable performance in diverse settings. Performance metrics may fail to capture clinically relevant aspects or models’ practical usability in clinical settings, thus concealing their limited clinical utility. As a result, such models may increase operating times, introduce workflow inefficiencies, or place additional cognitive burden on clinicians. |
| Selection of sub-optimal algorithms | Selection of sub-optimal algorithms refers to the risk of choosing models based on misleading or incomplete validation rather than true clinical utility. If performance metrics do not accurately reflect real-world applicability, models may be selected that perform well in development settings but lead to errors, inefficiencies, or reduced quality of care in clinical practice. Conversely, more effective algorithms may be overlooked, limiting clinical benefit and slowing progress in AI-assisted surgery. |
| Compromised clinical safety | Compromised clinical safety refers to the potential of surgical AI systems to contribute to medical errors, adverse events, or unsafe clinical decisions. This may occur when models produce inaccurate or misleading outputs, are used beyond their validated scope, are applied in situations in which their behavior is not well understood, or from flawed validation practices, such as inappropriate metric use, data leakage, or poorly designed validation, that fail to reveal limitations prior to clinical use. In such cases, AI use can lead to delays, complications, or inappropriate interventions that put patient safety at risk. |
| Loss of trust in surgical AI | Loss of trust in surgical AI refers to clinicians and other stakeholders becoming hesitant to adopt or rely on AI models when these systems fail to perform reliably or align with expectations. This may be caused by repeated failures, biases, poor generalization, unpredictable behavior, or lack of transparency. When models fail to match expectations set during development or produce unreliable predictions, unexplained errors, or inconsistencies across different clinical settings, confidence in their usefulness declines, potentially also reducing clinicians’ willingness to engage with or contribute to future surgical AI research and development. Such loss of trust may also affect regulators, institutional stakeholders, and investors, and can raise concerns around ethical responsibility and accountability in the use of surgical AI. |
| Stagnation of research progress | Stagnation of research progress refers to a slowdown in innovation and advancement within surgical AI due to unreliable validation practices or non-comparable benchmarks. When flawed metrics, poor generalization, or inconsistent reporting, among others, obscure the true algorithm performance, it becomes difficult to identify meaningful improvements or reproduce prior findings. As a result, researchers may repeatedly explore already-solved problems, waste resources on uninformative comparisons, or fail to translate experimental insights into clinical progress, ultimately hindering scientific development in the field. |
| Regulatory rejection or delay | Regulatory rejection or delay refers to the failure of surgical AI systems to obtain timely approval for clinical use due to insufficient or unconvincing validation or compromised quality control processes. Regulatory pathways are intended to safeguard patient safety and public trust. When AI systems do not adequately demonstrate clinical benefit, fairness, safety, effectiveness, generalizability, or reliability, approval may be withheld – delaying or preventing their availability for real-world use. |
| Investment risk | Investment risk refers to the potential for financial, institutional, or strategic resources to be committed to surgical AI systems that ultimately fail to deliver clinical value or broader impact. This may result from flawed validation, misleading performance claims, or inadequate alignment with real-world needs. As a result, organizations may experience financial losses, reputational damage, or delays in innovation. |
| Lack of clinical/real-world adoption of surgical AI | Lack of clinical adoption of surgical AI refers to the failure of AI models to be integrated in real-world settings. Challenges such as misleading performance metrics, poor generalization, usability issues, or a mismatch between expected and actual clinical performance can create barriers to the integration and acceptance of surgical AI in the clinical workflow. Even if a model shows strong results in experimental settings, a lack of validation in real-world conditions - such as different surgical teams, workflows, or patient populations - can limit its adoption. If AI tools do not align with clinical needs or fail to provide tangible improvements over existing methods, they may struggle to gain acceptance and practical use in surgical settings. |
| Reduced or delayed patient/caregiver benefit | Reduced or delayed patient and caregiver benefit refers to the diminished impact or delayed realization of benefits of surgical AI on patient outcomes and clinical support when models fail to generalize, align with clinical needs, or perform reliably in practice. This may follow from performance expectations that were not met in real-world settings. Instead of enhancing surgical precision, efficiency, or safety, unreliable models may introduce errors, delays, or additional workload for healthcare professionals. In the worst cases, AI-driven mistakes can lead to complications or adverse events, ultimately harming patients rather than improving care. |
| Negative environmental impact | Negative environmental impact refers to the unintended ecological consequences of inefficient or unvalidated surgical AI pipelines. Poorly designed benchmarks and redundant training of sub-optimal models can lead to excessive computational resource use, unnecessary data processing, and inflated carbon emissions. Moreover, lack of reproducibility or transparency may cause repeated experiments or model retraining without added value. Together, these factors contribute to an avoidable environmental burden that undermines the sustainability of surgical AI research and deployment. |
Appendix SUPPL. NOTE 3 Results of the systematic review per pitfall
While pitfalls can theoretically occur in any validation study, their actual prevalence in state-of-the-art surgical AI publications remained unclear. To address this, we conducted a systematic screening of all papers at the 2023 Medical Image Computing and Computer Assisted Intervention (MICCAI) conference that applied deep learning methods to surgical data. In this section, we present the results for each of the identified pitfalls.
[P1] Pitfalls related to data
P1.1: Non-representativeness, low quality, and low relevance of data
Was the proposed model tested on out-of-distribution data (e.g., data from different centers or different surgeries)?
Yes: 19.6%
No: 54.4%
Unclear: 26.1%
Were parts of the data excluded? (n = 46)
Yes: 17.4% (with clear criteria: 15.2%; without clear criteria: 2.2%)
No: 41.3%
Unclear: 41.3%
Were the datasets public or private? (n = 46)
Public datasets: 41.3%
Private datasets: 32.6%
Combination of private and public datasets: 19.6%
New dataset(s) to be made public: 4.4%
Unclear: 2.2%
Which datasets were used?
63% of tasks specified the exact datasets used (n = 29)
A total of 38 datasets were used
79% of datasets were only used once
The data sets used the most were EndoVis 2018 (used by 17.2% of tasks), Cholec80 (13.8%) and EndoVis 2017 (10.3%)
P1.2: Limited sample size/test cohorts, or data imbalance
Number of videos
| Minimum | Mean | Median | Maximum | |
| Training | 6 | 234.2 | 37 | 1,500 |
| Validation | 2 | 111.4 | 10 | 1,077 |
| Test | 2 | 141.6 | 22 | 1,321 |
Were data set sizes reported? (n = 46)
Yes: 41.3%
Partially: 30.4%
No: 21.7%
Unclear: 6.5%
Was the number of classes within the dataset reported? (n = 39)
Yes, they reported in detail for the general data: 48.7%
Yes, they reported in detail for each of the data subsets: 2.6%
Partially: 10.3%
No: 35.9%
Unclear: 2.6%
Was reported how classes are distributed in the dataset? (e.g., the percentage of each class in each video) (n = 39)
Yes, they reported in detail for the general data: 5.1%
Yes, they reported in detail for each of the data subsets: 2.6%
Partially: 7.7%
No: 84.6%
P1.3: Existence of spurious correlations within data
Were spurious correlations considered by the authors? (n = 46)
Considered in models: 2.2%
Mentioned by authors, not considered: 6.5%
No: 89.1%
Unclear: 2.2%
P1.4: Incomplete annotation or missing contextual information
Were all frames in the test set(s) annotated? (n = 42)
Yes: 48.8%
No: 17.0%
Unclear: 34.2%
P1.5: Unreliable or inconsistent annotation
Were object instances tracked (manually) over time for validation? (n = 33)
Yes: 0.0%
No: 42.4%
Unclear: 57.6%
P1.6: Poor data splitting
Which principle best described data splitting ? (n = 46)
Train / validation / test: 31.3%
Train / test (single split, no validation): 27.1%
k-fold CV (train / validation) / test: 10.4%
k-fold CV (no test): 6.3%
Train / validation (single split, no test): 2.1%
k-fold CV (no test): 2.1%
Mixture: 2.1%
Other: 2.1%
No splitting: 2.1%
Unclear: 14.6%
What were potential sources of data leakage? (n = 45)
None: 26.7%
Hierarchies are not handled properly: 26.7%
Non-independence between training and test samples: 20.0%
Lack of clean separation of training and test dataset: 17.8%
No test set: 17.8%
Unclear: 17.8%
Pre-processing on training and test set: 4.4%
Feature selection on training and test set: 2.2%
Lack of description of data split: 2.2%
Model uses features that are not legitimate: 2.2%
No concrete information on data split: 2.2%
Not enough info on which model is trained with which dataset: 2.2%
Temporal leakage: 2.2%
Was the test set untouched? (n = 43)
Yes: 46.5%
No: 14.0%
Other: 2.3%
Unclear: 37.2%
[P2] Pitfalls related to metric selection and configuration
P2.1: Mismatch of metrics and clinical needs
Were metric choices justified? (n = 46)
Yes: 30.6% (by popularity: 20.4%)
No: 69.4%
Were clinical needs reflected in the metric choice? (n = 46)
Yes: 6.5%
Partially: 4.4%
No: 8.7%
Unclear: 80.4%
P2.2: Lack of common metrics that assess temporal aspects
Were algorithm properties specific to temporal data assessed? (n = 35)
Yes: 8.6%
Partially: 8.6%
No: 77.2%
Unclear: 5.7%
1 paper introduced a temporal consistency metric.
P2.3: Inappropriate metric selection for handling annotation uncertainties
Were event boundaries handled in a specific way? (only for process-focus tasks; n = 11)
Yes: 0.0%
No: 81.8%
Unclear: 18.2%
P2.4: Lack of metric robustness across varying real-world conditions
Was the frame rate considered in the validation? (n = 34)
Yes: 17.7%
Partially: 5.9%
No: 58.8%
Unclear: 17.7%
Was the image quality considered in the validation (e.g., resolution or complexity)? (n = 46)
Yes: 8.7%
Partially: 6.5%
No: 56.5%
Unclear: 28.3%
P2.5: Non-standardized configuration and definition of metrics
Were non-standardized metrics considered for model validation? (n = 46)
Yes: 26.1%
No: 67.4%
Unclear: 6.5%
Was explicitly described how the metric was computed (including hyperparameter choice)? (n = 46)
Yes: 8.7%
Partially: 6.5%
No: 82.6%
Unclear: 2.2%
P2.6: Non-suitability of hyperparameters from unrelated domains
Did the metrics contain hyperparameters? (n = 46)
Yes: 32.6%
No: 47.8%
Unclear: 19.6%
If so, were hyperparameters justified? (n = 23)
Yes: 8.7% (by popularity: 4.4%)
No: 56.5%
Unclear: 30.4%
[P3] Pitfalls related to metric aggregation and reporting
P3.1: Non-independence within the test set
Were test cases independent? (n = 46)
Yes: 10.9%
No: 28.3%
Unclear: 58.7%
Other: 2.2%
If hierarchies were present in the data, were they properly addressed? (n = 39)
Yes: 5.1%
Partially: 5.1%
No: 23.1%
Unclear: 66.7%
Was the video length considered in the validation? (n = 34)
Yes: 14.7%
Partially: 2.9%
No: 58.8%
Unclear: 23.5%
P3.2: Clinically uninformative aggregation
Was relevance considered in aggregation? (n = 46)
Yes: 0.0%
No: 34.8%
Unclear: 65.2%
Were results stratified with respect to relevant aspects (e.g., object size / shape, sensor quality, …)? (n = 46)
Yes: 28.3%
Partially: 2.2%
No: 69.6%
P3.3: Lack of contextualization of performance values
Were performance values put into context (e.g., by including inter-rater agreement, by defining of what constitutes a meaningful difference, or by defining of what constitutes a value sufficient to solve the underlying clinical task)? (n = 46)
Yes: 13.0%
Partially: 6.5%
No: 80.4%
P3.4: Lack of uncertainty reporting
Manner of reporting variability or uncertainty (n = 46)
None: 39.1%
Standard deviation: 23.9%
Values with +/- (unclear whether it is standard deviation): 21.7%
Interquartile Range (graph): 17.4%
Graphs (other): 4.4%
Confidence intervals: 2.2%
Prediction intervals: 2.2%
Standard deviation in graphs: 2.2%
Standard error: 2.2%
Variability of different runs: 2.2%
Statistical tests: 2.2%
If standard deviation was reported, how was it calculated? (n = 21)
Not reported: 81.0%
From cross-validation (over non-overlapping samples): 4.8%
From cross-validation (unclear): 4.8%
Over different runs: 4.8%
Other: 4.8%
P3.5: Insufficient reporting
Was the reporting comprehensive? (n = 46)
Comprehensive reporting: 2.2%
Only partly described: 97.8%
1 single paper mentioned a reporting guideline but did not fully follow it.
Was the aggregation procedure described in detail? (n = 41)
Yes: 4.9%
Partially: 29.3%
No: 56.1%
Unclear: 9.8%
Ethical, Legal, Societal Aspects (ELSA):
Were ethical aspects reported? (n = 46)
Yes: 15.2%
Partially: 6.5%
No: 78.3%
Were fairness / bias / equity aspects reported? (n = 46)
Yes: 4.4%
Partially: 6.5%
No: 89.1%
Were social / legal / governance aspects reported? (n = 46)
Yes: 2.2%
Partially: 6.5%
No: 91.3%